Juhasz Corporation makes a product with the following standards for direct labor and variable overhead:

Marvel Parts, Inc., manufactures auto accessories. One of the company’s products is a set of seat covers that can be adjusted to fit nearly any small car. The company has a standard cost system in use for all of its products. According to the standards that have been set for the seat covers, the factory should work 990 hours each month to produce 1,980 sets of covers. The standard costs associated with this level of production are:

Total Per Set
of Covers
Direct materials $ 39,798 $ 20.10
Direct labor $ 5,940 3.00
Variable manufacturing overhead (based on direct labor-hours) $ 3,168 1.60
$ 24.70

During August, the factory worked only 1,000 direct labor-hours and produced 2,200 sets of covers. The following actual costs were recorded during the month:

Total Per Set
of Covers
Direct materials (7,400 yards) $ 40,700 $ 18.50
Direct labor $ 8,140 3.70
Variable manufacturing overhead $ 3,960 1.80
$ 24.00

At standard, each set of covers should require 3.0 yards of material. All of the materials purchased during the month were used in production.

Required:

1. Compute the materials price and quantity variances for August.

2. Compute the labor rate and efficiency variances for August.

3. Compute the variable overhead rate and efficiency variances for August.

(Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)

In August the company produced 7,900 units using 4,080 direct labor-hours. The actual variable overhead cost was $15,096. The company applies variable overhead on the basis of direct labor-hours.

The variable overhead rate variance for August is:

Multiple Choice

  • $1,185 F

  • $1,224 U

  • $1,185 U

  • $1,224 F
Answer #1

Material Price Variance = (Standard Price – Actual Price)*Actual Quantity Purchased

= (20.10/3 – Actual Rate)*7400

= 6.7*7400 – 40,700

= $8,880 Favorable

Material Quantity Variance = (Standard Quantity – Actual Quantity Used)*Standard Price

= (2200*3 – 7400)*6.7

= $5,360 Unfavorable

Labor Rate Variance = (Standard Rate – Actual Rate)*Actual Hours

= (3*2 – Actual Rate)*1000

= $2,140 Unfavorable

Efficiency Variance = (standard hours – actual hours)*Standard Rate

= (2200/2 – 1000)*6

= $600 favorable

Variable overhead rate variance = (Standard Rate – Actual Rate)*Actual Hours

= (1.6*2 – Actual Rate)*1000

= $760 U

Efficiency Variance = (standard hours – actual hours)*Standard Rate

= (100)*3.2

= $3200 F